# Compositional simulation-based inference for time series

### v1

@@ 3.2.2 Local score composition
At inference time, we have to approximate $s_{\text{glob}}({\bm{\theta}}_{a},{\bm{x}}_{o}^{0:T})=\nabla_{{\bm{\theta}}_{a}}\log p({\bm{\theta}}_{a}|{\bm{x}}_{o}^{0:T})$, by composing all local scores estimated by the network (Algorithm 2 the compose function). In the end, $s_{\text{glob}}$ will be passed to a diffusion sampler to obtain samples from the global posterior $p({\bm{\theta}}|{\bm{x}}_{o}^{0:T})$.
[…]
FNPE: Geffner et al. (2023) avoids the issue by constructing an alternative sequence of distributions $q_{a}~{}(0\leq a\leq A)$ that satisfies $q_{0}=p({\bm{\theta}}|{\bm{x}}^{0:T})$ and $q_{A}=\mathcal{N}({\bm{\theta}};\mu_{A},\frac{\sigma_{A}^{2}}{T}I)$. Adapting their proposal to the Markovian case, we obtain
While this approach is generally applicable, it renders the backward SDE (Eq. 2) invalid as the marginals no longer follow the corresponding forward diffusion process. This requires the use of MCMC-based sampling corrections, such as unadjusted Langevin dynamics (Geffner et al., 2023) or similar approaches (Sjöberg et al., 2024).
GAUSS/JAC: Linhart et al. (2024) have recently proposed an alternative approach by approximating the marginal scores derived from the forward diffusion process. This enables the utilization of the reverse SDE and consequently allows for the application of standard diffusion-based sampling techniques. A second-order approximation (i.e. using Gaussian assumptions), suggests the approximation
where $\mathbf{\Lambda}({\bm{\theta}}_{a})=\sum_{t=0}^{T-1}\mathbf{\Sigma}_{a,t,t+1}^{-1}+(1-T)\mathbf{\Sigma}_{a}^{-1}$. This approach requires specifying two ”hyperparameters”: the denoising prior and posterior precision matrices. The prior precision matrix is typically known analytically, along with the marginal prior score. In contrast, the denoising posterior precision matrix must be estimated. To address this, GAUSS assumes a Gaussian posterior and analytically computes the denoising posterior covariances. Alternatively, JAC iteratively estimates it using Tweedie’s Moment projection (Boys et al., 2024), leveraging the Jacobian of the score network. Furthermore, the above approximation is only valid if $\mathbf{\Lambda}$ is positive definite. Unlike Linhart et al. (2024), who clip the diffusion process to a specific prior range to avoid numerical issues, we adapt all estimated posterior precision matrices to ensure positive definiteness of $\mathbf{\Lambda}$. This adaptation resolves the problem even for long time series (i.e., $T>100$; see Appendix C.2 for details). Unless otherwise specified, we use the GAUSS approximation as the default.
@@ 3.3 Factorized likelihood(-ratio) estimation (FNLE, FNRE)
For likelihood(-ratio) estimation, applying the approach outlined above is straightforward: by assumption, the likelihood factorizes as shown in Eq. 1. Consequently, we can instead directly learn the transition density $p({\bm{x}}^{t+1}|{\bm{x}}^{t},{\bm{\theta}})$, similarly to the method described in Alg. 1, by replacing the score matching loss with the appropriate likelihood (or likelihood-ratio) loss (Papamakarios et al., 2019; Hermans et al., 2020; 2022; Miller et al., 2022).
Once the transition density is obtained, the global log-likelihood approximation $\ell_{\text{glob}}$ can be computed directly from Eq. 1 (i.e., by simply summing up the local approximations). Classical MCMC techniques can be employed for sampling. We use reference implementations of NLE and NRE as implemented in the sbi package  (Tejero-Cantero et al., 2020b), adapted to the Markovian setting.
@@ 4.1 Evaluation approach
We empirically assess the accuracy and computational efficiency of the proposed approach, comparing it against non-factorized NPE with an appropriately chosen embedding net as a baseline. We note that the simulation budget is not determined by the number of simulator calls $N$, but rather by the number of calls to the state-transition function, i.e., $T\cdot N$. We, therefore, configure the NPE baseline with a $T_{\text{max}}=10$ steps, i.e., if the simulation budget is $10k$, the NPE baseline used $1000$ simulations, each of which is simulated for $10$ steps.We additionally use data augmentation by duplicating shortened variants ($T<T_{\text{max}}$) of these simulations. This is to help the RNN embedding net to generalize to different sequence lengths. To also amortize over different initial conditions, the initial condition for each ten-step simulation was drawn from the proposal also used in the factorized methods.
As main metrics for comparison, we use sliced Wasserstein distance (sW1) and Classifier two sample test accuracy (C2ST) on reference posterior samples (Bischoff et al., 2024; Lopez-Paz & Oquab, 2016; Kolouri et al., 2019). We average the estimated value over a total of $10$ randomly drawn observations. The whole process, i.e., training, sampling, and evaluation, was repeated five times starting from different random seeds.
We begin by evaluating the methods on a set of newly designed benchmark tasks for Markovian simulators (Sec. 4.2). Next, we apply the approach to classical models from ecology and epidemiology, including the stochastic Lotka-Volterra and SIR models (Sec. 4.3). Finally, we demonstrate the scalability of the method on a large-scale Kolmogorov flow task, where we perform inference on very high-dimensional data using only 200k simulator steps (Sec. 4.4)
@@ 4.2 Benchmarks
To investigate several properties of the proposed approach, we develop several synthetic tasks of first-order Markovian time series with associated reference posterior samplers (details in Appendix B.1).
Gaussian RW: A Gaussian Random Walk of form ${\bm{x}}^{t+1}=\alpha\cdot{\bm{x}}^{t}+{\bm{\theta}}+\epsilon$ for $\epsilon\sim\mathcal{N}(0,{\bm{I}}),\alpha<1.$ with ${\bm{x}}^{t}\in\mathbb{R}^{d}$ and ${\bm{\theta}}\in\mathbb{R}^{d}$. This simple task offers an analytic Gaussian posterior.
@@
Periodic/Linear SDE: Both tasks can be formulated as linear stochastic differential equation $d{\bm{x}}^{t}={\bm{A}}({\bm{\theta}}){\bm{x}}^{t}+{\bm{B}}({\bm{\theta}})d{\bm{w}}^{t}$. The periodic SDE is nonlinearly parameterized ${\bm{A}}$ to be oscillatory with ${\bm{\theta}},{\bm{x}}^{t}\in\mathbb{R}^{2}$. This task has a posterior with four modes symmetric around the origin. The linear SDE task s every entry of ${\bm{A}}$ and ${\bm{B}}$ with ${\bm{x}}^{t}\in\mathbb{R}^{3}$ and ${\bm{\theta}}\in\mathbb{R}^{18}$.
@@
We first examine the scalability of all evaluated methods as a function of the length and dimensionality of the time series, while maintaining a fixed simulation budget. Specifically, we assess each method on the Gaussian random walk (RW) example in 1, 2, and 10 dimensions, using a total of 10k simulations (Fig. 2a). Both FNLE and FNSE demonstrate superior simulation efficiency compared to the NPE baselines. FNRE, however, appears to be more sensitive to local errors, with these errors accumulating over time, resulting in a consistent decline in performance for long time series. A similar observation and hypothesis was made by Geffner et al. (2023) for i.i.d. data. Interestingly, this effect is largely reduced in the FNLE approach, which was not part of their empirical analysis (but successfully applied by Boelts et al. (2022)). On the other hand, the computational cost escalates significantly on long time series, as evaluating the global likelihood requires $T$ forward passes for each iteration of MCMC sampling, making FNLE and FNRE relatively slow (especially NLE, as a forward pass through the normalizing flow in FNLE is more costly than the classifier used in FNRE, Fig. 2a). In contrast, the sampling cost for FSNE remains moderate due to its more efficient sampling method, while performance remains similar to or better than FNLE.
@@
In contrast to FNLE and FNRE, FNSE can introduce errors when composing the local scores. Therefore, we investigate the scaling behavior of different score-composition techniques for long time series (Fig. 2b, Appendix Fig. 5b). Previous work within the i.i.d case did not investigate performance beyond thirty samples (Geffner et al., 2023; Linhart et al., 2024). Specifically, for time series exceeding a length of $T>100$, the FNPE method starts to become numerically unstable due to diverging trajectories. Although the assumptions of the GAUSS/JAC approximations are violated (as the Mixture RW has a non-Gaussian posterior), these methods still tend to perform well in practice in our implementation. Even when Langevin corrections (Song et al., 2021; Geffner et al., 2023) are introduced, the GAUSS and JAC method remain superior, especially for large $T$ (Fig. A7, Appendix Sec. C.2).
@@
Next, we conducted benchmarks across several tasks. Overall, FNSE and FNLE show higher accuracy compared to FNRE and NPE in most tasks (Fig. 2c, Appendix Fig. 5c). An exception is the double-well task, where NPE performs relatively well due to the SDE’s rapid convergence to a stationary distribution, enabling the RNN to learn a generalizable summary statistic for inference. FNSE strongly outperforms other methods in cases of multimodal transition densities (e.g., Mixture RW) and performs comparably to likelihood-based approaches in simpler cases (e.g., Periodic/Linear SDE) even if the posterior is strongly multimodal (Periodic SDE). Learning from single-step transitions is an extreme end of a spectrum of possibilities. We therefore also evaluate this benchmark on partially factorized variants using five transitions each (Appendix Fig. 6). Overall, this could improve (e.g., Double Well), but it could also hurt (e.g., Mixture RW) performance relative to the NPE baseline.
@@
Finally, we investigate the impact of the proposal distribution, specifically for the FNSE method trained over 10k simulation steps (Fig. 2d, extended results in Fig. A8). We use a Normal distribution centered at zero, adjusting its standard deviation ranging from $\sigma=0.1$ (too narrow) to $\sigma=1000$ (too wide). Our results indicate a relatively broad range of standard deviations that result in good performance. If the distribution is too narrow, the time series observation may reach values outside the training domain, while if it is too wide, it the model is trained on values not observed during evaluation.
@@
Fig 3: Lotka Volterra and SIR experiments : The FNSE approximate posterior (predictive) of the best performing FNSE model using 100k transition steps to train, visualized on subsequences from a fixed observation and associated true parameter. (a, b). We then show the quantitative performance in terms of C2ST and sW1 for each task on ten randomly selected observations, each run repeated five times (c, d).
@@ 4.3 Lotka Volterra and SIR
To further evaluate our method, we tested it on two famous models from ecology and epidemiology: the Lotka-Volterra and Susceptible-Infected-Recovered (SIR) simulators. The Lotka-Volterra simulator models predator-prey dynamics through four parameters that govern prey and predator growth, hunting rates, and mortality. Populations evolve over time according to interconnected stochastic differential equations. The SIR (Susceptible-Infected-Recovered) simulator is a fundamental model for understanding infectious disease spread. Although commonly used for benchmarking SBI on fixed-size observations Lueckmann et al. (2021), we adapted these models for our analysis (details in Appendix Sec. B.1).
@@
We trained all methods on this task using $100k$ transition simulations. We then evaluate the accuracy of all models for performing inference on sequences of length $T=1,10,100$. For both tasks we see that the factorized methods (except FNRE; in LV) generally outperform the NPE baseline (Fig. 3cd). NFSE, in particular, significantly outperforms any other approaches on the SIR task.

### v2

@@ 3.2.2 Local score composition
At inference time, we need to recover the global score $s_{\text{glob}}({\bm{\theta}}_{a}|{\bm{x}}_{o}^{0:T})=\nabla_{{\bm{\theta}}_{a}}\log p({\bm{\theta}}_{a}|{\bm{x}}_{o}^{0:T})$, by composing all local scores $s_{\text{local}}({\bm{\theta}}_{a}|{\bm{x}}_{o}^{t:t+1})=\nabla_{{\bm{\theta}}_{a}}\log p({\bm{\theta}}_{a}|{\bm{x}}_{o}^{t:t+1})$ estimated by the network, thereby defining the compose function in Alg. 2. Finally, the composed global score $s_{\text{glob}}$ will be passed to a diffusion sampler to obtain samples from the target posterior $p({\bm{\theta}}|{\bm{x}}_{o}^{0:T})$.
While the composition as introduced in Eq. 3 is valid for $a=0$, it becomes invalid for any $a>0$ (i.e., each noisy posterior). The reason for this is the following: The likelihood for the ‘noisy’ parameter $p({\bm{x}}^{0:T}|{\bm{\theta}}_{a})$ no longer satisfies the Markov property, even if $p({\bm{x}}^{0:T}|{\bm{\theta}})$ does (Weilbach et al., 2023; Gloeckler et al., 2024b; Geffner et al., 2023; Rozet & Louppe, 2023). In the i.i.d. setting, this issue has been tackled by Linhart et al. (2024), and we will adapt their approach to the Markov setting. The correct global score is intractable in most cases, but they proposed an approximation that can be directly adapted for Markovian time series:
$\nabla_{{\bm{\theta}}_{a}}\log p({\bm{\theta}}_{a}|{\bm{x}}^{0:T})\approx\mathbf{\Lambda}({\bm{\theta}}_{a})^{-1}\left(\sum_{t=0}^{T-1}\mathbf{\Sigma}_{a,t,t+1}^{-1}s_{\phi}({\bm{\theta}}_{a}|{\bm{x}}^{t},{\bm{x}}^{t+1})+(1-T)\mathbf{\Sigma}_{a}^{-1}\nabla_{{\bm{\theta}}_{a}}\log p({\bm{\theta}}_{a})\right),$
where $\mathbf{\Lambda}({\bm{\theta}}_{a})=\sum_{t=0}^{T-1}\mathbf{\Sigma}_{a,t,t+1}^{-1}+(1-T)\mathbf{\Sigma}_{a}^{-1}$. Here $\mathbf{\Sigma}_{a}$ denotes the covariance matrix of $p({\bm{\theta}}|{\bm{\theta}}_{a})$ (the usually tractable denoising prior) and $\mathbf{\Sigma}_{a,t,t+1}$ denotes the covariance matrix of $p({\bm{\theta}}|{\bm{\theta}}_{a},{\bm{x}}^{t,t+1})$ (the denoising posterior), which we need to estimate. This can be done by estimating the posterior covariance from samples of the local posteriors obtained via the diffusion model, referred to as GAUSS. Unless otherwise specified, we use this approximation as the default composition rule for FNSE. Alternatively, we can estimate it via Tweedie’s moment projection using the Jacobian of the score estimator, referred to as JAC (Linhart et al. (2024), see Appendix B for details). In contrast, Geffner et al. (2023) addressed this issue (in the i.i.d. setting) through post-hoc sampling corrections (we referred to this uncorrected variant as FNPE, Appendix Sec. B.2 for details)
@@ 3.3 Factorized likelihood(-ratio) estimation (FNLE, FNRE)
For likelihood(-ratio) estimation, applying our approach is straightforward: by assumption, the likelihood factorizes as shown in Eq. 1. As a consequence, we can directly learn the transition density $p({\bm{x}}^{t+1}|{\bm{x}}^{t},{\bm{\theta}})$, similarly to the method described in Alg. 1, by replacing the score matching loss with the appropriate likelihood (or likelihood-ratio) loss (Papamakarios et al., 2019; Hermans et al., 2020; 2022; Miller et al., 2022). Once the transition density is obtained, the global log-likelihood approximation $\ell_{\text{glob}}$ can be computed directly from Eq. 1 (i.e., by simply summing up the local approximations). Classical MCMC techniques can be employed for sampling. We use reference implementations of NLE and NRE as implemented in the sbi package  (Tejero-Cantero et al., 2020b; Boelts et al., 2024), adapted to the Markovian setting.
3.4 Construction of the proposal distribution
A proposal $\tilde{p}({\bm{x}}^{t})$ for single state transitions needs to satisfy conditions (i) and (ii) in Sec. 3.2.1, but solely relying on these properties does not necessarily make up a “good” proposal. Essentially, it specifies which regions of the data domain are represented in the training dataset. Therefore, a better local posterior approximation is expected for states that are more likely to be generated by the proposal compared to those that are less likely. This provides an opportunity and challenge to design appropriate proposals for a given simulator.
For an amortized posterior approximation, the proposal distribution design should be guided by the prior predictive distribution. The resulting trajectories ${\bm{x}}^{0:T}\sim p({\bm{x}}^{0:T})$ encompass likely states ${\bm{x}}^{t}$ and we can use this to construct a reasonable proposal $\tilde{p}({\bm{x}}^{t})$ (e.g. by simply proposing these at random)—however, this approach comes at the cost of upfront simulations. Alternatively, we can construct the proposal $\tilde{p}({\bm{x}}^{t})$ based on observations available prior to training. This will focus simulation resources on the intended application but might not yield a good amortized approximation. For further details on proposal design, see Appendix Sec. C.
@@ 4.1 Evaluation approach
We empirically assess the accuracy and computational efficiency of the proposed approach, comparing it against non-factorized NPE with an appropriately chosen embedding net as a baseline. We note that the simulation budget is not determined by the number of simulator calls $N$, but rather by the number of calls to the state-transition function, i.e., $T\cdot N$. We, therefore, configure the NPE baseline with a $T_{\text{max}}=10$ steps, i.e., if the simulation budget is $10k$, the NPE baseline used $1000$ simulations, each of which is simulated for $10$ steps (experiments with longer segments in Appendix E). We additionally use data augmentation by duplicating shortened variants ($T<T_{\text{max}}$) of these simulations. This aims to help the RNN embedding net to generalize to different sequence lengths. To also amortize over different initial conditions, the initial condition for each ten-step simulation was drawn from the proposal also used in the factorized methods. As main metrics for comparison, we use sliced Wasserstein distance (sW1) and Classifier two sample test accuracy (C2ST) on reference posterior samples (Lopez-Paz & Oquab, 2016; Kolouri et al., 2019; Bischoff et al., 2024). We average the estimated value over a total of $10$ randomly drawn observations. The whole process, i.e., training, sampling, and evaluation, was repeated five times, starting from different random seeds.
We begin by evaluating the methods on a set of newly designed benchmark tasks for Markovian simulators (Sec. 4.2). Next, we apply the approach to classical models from ecology and epidemiology, including the stochastic Lotka-Volterra and SIR models (Sec. 4.3). Finally, we demonstrate the scalability of the method on a large-scale Kolmogorov flow task, where we perform inference on very high-dimensional data using only 200k simulator steps (Sec. 4.4).
@@ 4.2 Benchmarks
To investigate several properties of the proposed approach, we develop several synthetic tasks of first-order Markovian time series with associated reference posterior samplers (details in Appendix D.1).
Gaussian RW: A Gaussian Random Walk of form ${\bm{x}}^{t+1}=\alpha\cdot{\bm{x}}^{t}+{\bm{\theta}}+\epsilon$ for $\epsilon\sim\mathcal{N}(0,{\bm{I}}),\alpha<1$ with ${\bm{x}}^{t}\in\mathbb{R}^{d}$ and ${\bm{\theta}}\in\mathbb{R}^{d}$. This simple task offers an analytic Gaussian posterior.
@@
Periodic/Linear SDE: Both tasks are governed by the linear SDE $d{\bm{x}}^{t}={\bm{A}}({\bm{\theta}}){\bm{x}}^{t}+{\bm{B}}({\bm{\theta}})d{\bm{w}}^{t}$. The periodic SDE is oscillatory with ${\bm{\theta}},{\bm{x}}^{t}\in\mathbb{R}^{2}$ and a posterior with four modes. The linear SDE involves ${\bm{A}}$ and ${\bm{B}}$ with ${\bm{x}}^{t}\in\mathbb{R}^{3}$ and ${\bm{\theta}}\in\mathbb{R}^{18}$.
@@
We first examine the scalability of all evaluated methods as a function of the length and dimensionality of the time series while maintaining a fixed simulation budget. Specifically, we assess each method on the Gaussian random walk (RW) example in 1, 2, and 10 dimensions, using a total of 10k simulations (Fig. 2a). Both FNLE and FNSE show better simulation efficiency than the NPE baselines. FNRE is more sensitive to local errors, leading to a decline in performance over a long time series. A similar observation was made by Geffner et al. (2023) for i.i.d. data. Notably, this issue is reduced in the FNLE approach, which was not analyzed in their work but successfully applied by Boelts et al. (2022). On the other hand, the computational cost escalates significantly on long time series, as evaluating the global likelihood requires $T$ forward passes for each iteration of MCMC sampling, making FNLE and FNRE relatively slow (especially NLE, as a forward pass through the normalizing flow in FNLE is more costly than the classifier used in FNRE, Fig. 2a). In contrast, the sampling cost for FSNE remains moderate due to its more efficient sampling method, while performance remains similar to or better than FNLE.
@@
In contrast to FNLE and FNRE, the approximative score composition (for $a>0$) might impact the sampling quality for FNSE. Therefore, we investigate the scaling behavior of different score composition techniques for long time series (Fig. 2b, Appendix Fig. 8b). Previous work within the i.i.d. case did not investigate performance beyond thirty samples (Geffner et al., 2023; Linhart et al., 2024). Specifically, for time series exceeding a length of $T>100$, the FNPE method starts to become numerically unstable due to diverging trajectories. Although the assumptions of the GAUSS/JAC approximations are violated (as the Mixture RW has a non-Gaussian posterior), these methods still tend to perform well in practice in our implementation. Even when Langevin corrections (Song et al., 2021; Geffner et al., 2023) are introduced, the GAUSS and JAC method remain superior, especially for large $T$ (Fig. A10, Appendix Sec. B.2).
@@
Next, we conducted benchmarks across several tasks. Overall, FNSE and FNLE show higher accuracy compared to FNRE and NPE in most tasks (Fig. 2c, Appendix Fig. 8c). An exception is the double-well task, where NPE performs relatively well due to the SDE’s rapid convergence to a stationary distribution, enabling the RNN to learn a generalizable summary statistic for inference. FNSE strongly outperforms other methods in cases of multimodal transition densities (e.g., Mixture RW) and performs comparably to likelihood-based approaches in simpler cases (e.g., Periodic/Linear SDE) even if the posterior is strongly multimodal (Periodic SDE). Learning from single-step transitions is the extreme end of a spectrum of possibilities. We, therefore, also evaluate this benchmark on partially factorized variants using five transitions each (Appendix Fig. 9). Overall, this could improve (e.g., Double Well) or hurt (e.g., Mixture RW) performance relative to the NPE baseline. We also compared method to a broader spectrum of applicable baselines (Chen et al. (2021; 2023); Sharrock et al. (2024), Appendix Table A3, A4), and plot our results also against training budget (Appendix Fig. A6).
@@
Finally, we investigate the impact of the proposal distribution, specifically for the FNSE method trained over 10k simulation steps (Fig. 2d, extended results in Fig. A11). We use a Normal distribution centered at zero, adjusting its standard deviation ranging from $\sigma=0.1$ (too narrow) to $\sigma=1000$ (too wide). Our results indicate a relatively broad range of standard deviations of the proposal that result in good performance. If the distribution is too narrow, the observation may reach values outside the training domain, while if it is too wide, the model is trained on values not observed during evaluation. In addition, we compared our default choice to proposals constructed using prior predictive samples (using a subset of the simulation budget, Sec. 3.4, Appendix Sec. E.4). Overall, this leads to similar results even if the simulation budget used to construct the proposal is subtracted from the training budget (Tab. 1)
@@
Fig 3: Lotka Volterra and SIR experiments : The FNSE approximate posterior (predictive) of the best performing FNSE model using 100k transition steps to train, visualized on subsequences from a fixed observation and associated true parameter. (a, b). We then show the quantitative performance in terms of C2ST and sW1 for each task on ten randomly selected observations; each run is repeated five times (c, d).
@@ 4.3 Lotka Volterra and SIR
To further evaluate our method, we tested it on two famous models from ecology and epidemiology: the Lotka-Volterra and Susceptible-Infected-Recovered (SIR) simulators. The Lotka-Volterra simulator models predator-prey dynamics through four parameters that govern prey and predator growth, hunting rates, and mortality. The SIR simulator is a fundamental model for understanding infectious disease spread. Although commonly used for benchmarking SBI on fixed-size observations Lueckmann et al. (2021), we adapted these models for our analysis (details in Appendix Sec. D.1).
@@
We trained all methods on this task using $100k$ transition simulations. We then evaluate the accuracy of all models for performing inference on sequences of length $T=1,10,100$. For both tasks, we see that the factorized methods (except FNRE in Lotka Volterra) generally outperform the NPE baseline (Fig. 3cd, also Appendix Fig. 6 per training budget, Appendix Tab. 3 for additional baselines). NFSE, in particular, significantly outperforms all other approaches on the SIR task. We additionally performed a simulation-based calibration analysis (Talts et al., 2018) for FNSE and NPE, which, in line with these results, showed that FNSE is better calibrated than NPE (Fig. A7.). Notably, on the SIR task, we found NPE—in contrast to FNSE—to be unable to generalize to longer time series. In addition, we compared our default proposal to one constructed using simulations. For the Lotka Volterra task, this constructed proposal performed significantly better than our simple default (Appendix Tab. 2, Fig. 5).